# ToolACE-MT: Non-Autoregressive Generation for Agentic Multi-Turn Interaction

### v1

@@ Abstract
Agentic task-solving with Large Language Models (LLMs) requires multi-turn, multi-step interactions, often involving complex function calls and dynamic user-agent exchanges. Existing simulation-based data generation methods for such scenarios rely heavily on costly autoregressive interactions between multiple LLM agents, thereby limiting real-world performance of agentic tasks. In this paper, we propose a novel Non-Autoregressive Iterative Generation framework, called ToolACE-MT, for constructing high-quality multi-turn agentic dialogues. ToolACE-MT generates full conversational trajectories through three stages: coarse-grained initialization, iterative refinement, and offline verification. The initialization phase builds a structurally complete yet semantically coarse dialogue skeleton; the iterative refinement phase introduces realistic complexities and continued refinement via mask-and-fill operations; and the offline verification phase ensures correctness and coherence via rule- and model-based checks. Experiments demonstrate that ToolACE-MT enables efficient, effective and generalizable agentic data generation, offering a new paradigm for high-quality data construction in tool-augmented LLM scenarios.
Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable abilities in open-ended generation, reasoning, and instruction following (Guo et al. 2025; Wang et al. 2024; Jiang et al. 2024). Beyond passive language understanding, a growing frontier in LLM research involves agentic task-solving, where models take on the role of autonomous agents interacting with users and environments over multi-turn dialogues (Wang et al. 2023; Luo et al. 2025). These settings often involve multiple function calling111In this paper, function calling, tool calling and tool use are used interchangeably. and adaptive decision-making, significantly broadening the applicability of LLMs in real-world scenarios.
@@
To enable such agentic capabilities, high-quality multi-turn multi-step interaction data is essential. Multi-turn refers to multiple exchanges between the user and the assistant, while multi-step denotes task completion that requires executing a sequence of dependent actions, often through function calls. Together, they reflect the complexity of real-world agentic scenarios where task states are partially observable. However, constructing such data is inherently challenging: it requires generating complex but solvable tasks, maintaining coherent user-agent exchanges, and accurately simulating tool behaviors. A promising direction lies in multi-agent simulations, where multiple LLMs are assigned roles including user, assistant and tool to collaboratively generate full conversations through autoregressive interactions (Liu et al. 2025; Prabhakar et al. 2025). While effective at generating natural conversations, these approaches have several drawbacks: 1) They are computationally costly due to extended back-and-forth interactions, where each new turn must be generated one-by-one based on all previous context; 2) Task complexity and dialogue length are implicitly determined by model interactions and are difficult to constrain explicitly, which poses challenges for fine-grained data design; 3) Most critically, since assistant behaviors are generated autoregressively without access to global context, i.e. the overall task and dependencies between steps, it is difficult for the assistant to optimize the overall output structure and ensure consistency at each step. This lack of holistic awareness hinders factual accuracy, tool-use consistency, and task solvability, especially in scenarios requiring long-term planning. As a result, the quality of generated data largely depends on the capability of the LLMs playing the assistant role, often resembling a form of knowledge distillation from larger (assistant) models.
@@
In this paper, we propose ToolACE-MT, a novel framework for constructing multi-turn dialogues involving agentic tool-use behaviors, inspired by Non-Autoregressive Translation (NAT) and masked diffusion language models (Gu et al. 2018; Sahoo et al. 2024), which have been shown to be more efficient in language generation. Unlike traditional autoregressive multi-agent simulations (MAS), ToolACE-MT generates full conversational trajectories through a non-autoregressive pipeline consisting of three stages (see Figure 1):
Coarse-Grained Initialization: A coarse but structurally complete dialogue skeleton is generated, specifying overall tasks and action trajectory.
Iterative Refinement: Through carefully designed mask-and-fill procedures, the skeleton is progressively enriched with complexity injection and improved with reasonability refinement.
Offline Verification: Rule- and model-based checks are applied, filtering out inconsistent or invalid samples.
@@
Exerimental results on several agentic multi-turn benchmarks, including BFCL (Berkeley Function Calling Leaderboard) (Yan et al. 2024), -Bench (Yao et al. 2025) and ACEBench (Chen et al. 2025), show that models trained with ToolACE-MT generated data outperform those with autoregressive multi-agent simulation (MAS). Data analysis and ablation studies confirm the efficiency and effectiveness of our generation pipeline, and more experiments show the generalizability to different backbones.
@@
Related Work
Agentic Data Synthesis.
LLM agents equipped with external tools have shown realistic problem-solving capabilities (Qin et al. 2024; Gou et al. 2024; Lu et al. 2025). As current LLMs still face challenges with complex tasks (Mialon et al. 2024; Yao et al. 2025), learning from synthesized agentic data offers a promising direction. Early work focuses on single-turn synthesis, where the agent receives a one-time query and responds accordingly (Patil et al. 2023; Zeng et al. 2023; Qin et al. 2024; Liu et al. 2024). However, real-world tasks often involve multi-turn, multi-step interactions, prompting recent studies to explore conversational data synthesis via multi-agent simulations (Tang et al. 2023; Liu et al. 2025; Wang et al. 2025). Closest to our approach is Prabhakar et al. (2025), which adopts a two-stage synthesis process. While their first stage resembles ours which generates task configurations and ground-truth answers, they still rely on multi-agent simulations in the second stage to collect full interaction trajectories.
Agentic Model Training.
Fine-tuning on synthesized data remains central to agentic model training (Qin et al. 2024; Liu et al. 2025; Prabhakar et al. 2025). With reinforcement learning (RL) proving effective in enhancing LLM reasoning (Shao et al. 2024; Guo et al. 2025), agentic RL has emerged as a promising alternative for developing agentic capabilities (Feng et al. 2025; Qian et al. 2025; Zhang et al. 2025; Jin et al. 2025). While agentic RL may reduce the complexity of data synthesis by enabling learning from sparse or indirect supervision, high-quality synthesized data remains essential to guide and stabilize training.
Non-Autoregressive Generation.
To overcome the inefficiencies and quality limitations in certain scenarios of traditional autoregressive generation, where tokens are produced one-by-one, non-autoregressive approaches have been proposed (Gu et al. 2018; Xiao et al. 2023). These methods include CTC-based objectives (Libovický and Helcl 2018), iterative refinement with Mask-Predict (Ghazvininejad et al. 2019), and insertion and deletion strategies (Gu, Wang, and Zhao 2019). Inspired by this line of work, we extend the non-autoregressive paradigm to the turn level, enabling more efficient and coherent agentic dialogue data synthesis.
Method
Problem Formulation
@@
This interaction results in a sequence of alternating observations and actions, , where is the user’s initial message and is the tool output or user reply following . constitutes the target for data generation, i.e., multi-turn conversational data.
Overview
To construct high-quality conversational sequences , prior work applies multi-agent simulations (Liu et al. 2025; Prabhakar et al. 2025), where LLMs simulate the user (producing observations after assistant responses), the tools (producing outputs after function calls, can be either actual tools or simulated ones), and the assistant (producing actions). While this approach is shown to work, it is costly, hard to verify and complexity control. We propose a more efficient and controllable method: Non-Autoregressive Iterative Generation (named ToolACE-MT), inspired by non-autoregressive translation (NAT) and diffusion models (Gu et al. 2018; Sahoo et al. 2024).
Our generation pipeline consists of three main stages: 1) Initialization 2) Iterative Refinement, and 3) Offline Verification. Figure 2 illustrates the overall workflow. We describe each stage in detail below.
Coarse-Grained Initialization
@@
Task Initialization.
We begin by sampling or specifying a candidate tool list from a predefined tool pool (Liu et al. 2025; Wang et al. 2025). The overall task is then generated with the following components: 1) a set of subtasks (with predefined per instance), 2) the required tools for each subtask, 3) and the number of steps required for tool usage for each subtask.
This step serves as high-level planning, outlining the overall trajectory without finalizing all the details.
Trajectory Initialization.
@@
Iterative Refinement
In this stage, we enhance the initial trajectory through multiple refinement passes, improving both complexity and semantic coherence. Inspired by Masked-Predict (Ghazvininejad et al. 2019), we iteratively apply mask-and-extend or mask-and-fill to progressively improve the trajectory (see Figure 3).
Complexity Injection.
@@
Clarification turns: user providing incomplete information follows by assistant asking clarification questions.
Tool awareness: assistant recognizing unsupported tasks and user updating the tool list.
Error simulation: including tool call failures or instability, resulting assistant reflecting and adjusting actions.
@@
Offline Verification
Given the extensive use of LLMs in the aforementioned stages, hallucination remains a critical issue, particularly in long multi-turn dialogues involving large tool lists (Liu et al. 2023). To address this, we conduct offline verification on the refined trajectories using a hybrid approach that combines rule-based and model-based methods (Liu et al. 2024, 2025).
@@
Experiments
@@
Experimental Setup
Dataset Construction.
We construct in total 8000 training instances using ToolACE-MT for experiments. For comparison, we also construct 8000 instances with multi-agent simulation (MAS) method introduced in Wang et al. (2025). For fair comparison, we leverage the same candidate tool pool and LLM (GPT-4o-2024-11-20222https://chatgpt.com) for generation, the same offline verification is applied for both datasets.
Models.
We use LLaMA3.1-8B-Instruct (AI@Meta 2024) as the base model in our main experiments. Other models, including Qwen2.5-Instruct-series (Yang et al. 2024) (0.5B, 1.5B, 3B and 7B) and Qwen3-8B (Yang et al. 2025), are also tested to validate the generalizability of our method. To verify the effectiveness of our proposed three stages, we also train models with data without offline verification and iterative refinement for ablation study.
Benchmarks and Evaluation.
We conduct experiments on several representative benchmarks targeting on the multi-turn capabilities of tool-augmented LLMs, including the Berkeley Function Call Leaderboard (BFCL-v3) (Yan et al. 2024), -Bench (Yao et al. 2025), and ACEBench (Chen et al. 2025). As we focus on the realistic multi-turn capabilities, we mainly present and analyze results on the categories related to Multi-Turn categories. Results in single turn are also listed (for BFCL, while those for ACEBench listed in Appendix) to show the robustness.
Training Details.
Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA (Hu et al. 2022) for model training. All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of , following a cosine learning rate schedule with a warmup ratio of 0.1.
Main Results
Results on BFCL.
Table 1 shows the results on BFCL-v3. The results demonstrate that ToolACE-MT significantly improves multi-turn function calling accuracy, outperforming strong open-source and even some proprietary models (e.g. Gemini-2.5-Pro). Specifically, ToolACE-MT achieves a multi-turn accuracy, a absolute improvement over the base model Llama3.1-8B-Inst (), and even higher than models with larger sizes like Llama3.1-70B () and DeepSeek-V3 (). Compared to the Multi-Agent Simulation (MAS) (), ToolACE-MT also achieves consistently better results across all multi-turn subcategories. These findings highlight the effectiveness of our non-autoregressive data generation framework in constructing coherent, contextually grounded dialogues with accurate tool usage.
@@
Results on More Benchmarks.
| Models | Multi-Turn | Agent (EA) | Agent (PA) |
@@
The results on -Bench (shown in Table 3) show a consistent trend, with ToolACE-MT outperforming MAS baseline and the ablation models. Interestingly, the base model Llama3.1-8B-Inst obtains a higher score of in the Airline domain, surpassing all trained models. This counterintuitive outcome can be attributed to a known evaluation limitation in -Bench (Zhu et al. 2025): several instances define empty actions as the correct responses, assessing the assistant’s ability to recognize unsolvable user requirements. When a model lacks sufficient capability and consistently fails to produce valid function calls, it may coincidentally align with these empty actions and receive positive rewards, despite not demonstrating actual understanding. However, this phenomenon does not persist after training, which ultimately leads to lower evaluation scores.
| Models | -Retail | -Airline | Overall |
@@
Data Efficiency
Data Generation Efficiency and Model Choices.
| Method | Cost | Quality | Performance |
| MAS with GPT-4o | 275k | 61.1 | 64.17 |
| ToolACE-MT with GPT-4o | 188k | 72.3 | 65.41 |
| ToolACE-MT with GPT-4o-mini | 394k | 48.7 | 60.13 |

### v2

@@ Abstract
Agentic task-solving with Large Language Models (LLMs) requires multi-turn, multi-step interactions, often involving complex function calls and dynamic user-agent exchanges. Existing simulation-based data generation methods for such scenarios rely heavily on costly autoregressive interactions between multiple LLM agents, thereby limiting real-world performance of agentic tasks. In this paper, we propose ToolACE-MT, a novel Non-Autoregressive Iterative Generation framework for constructing high-quality multi-turn agentic dialogues. ToolACE-MT generates full conversational trajectories through three stages: coarse-grained initialization, iterative refinement, and offline verification. The initialization phase builds a structurally complete yet semantically coarse dialogue skeleton; the iterative refinement phase introduces realistic complexities and continued refinement via mask-and-fill operations; and the offline verification phase ensures correctness and coherence via rule- and model-based checks. Experiments demonstrate that ToolACE-MT enables efficient, effective and generalizable agentic data generation, offering a new paradigm for high-quality data construction in tool-augmented LLM scenarios.
1 Introduction
@@
Large Language Models (LLMs) have demonstrated remarkable abilities in open-ended generation, reasoning, and instruction following Guo et al. (2025); Wang et al. (2024); Jiang et al. (2024). Beyond passive language understanding, a growing frontier in LLM research involves agentic task-solving, where models take on the role of autonomous agents interacting with users and environments over multi-turn dialogues Wang et al. (2023); Luo et al. (2025). These settings often involve multiple function calling111In this paper, function calling, tool calling and tool use are used interchangeably. and adaptive decision-making, significantly broadening the applicability of LLMs in real-world scenarios.
@@
To enable such agentic capabilities, high-quality multi-turn multi-step interaction data is essential. Multi-turn refers to multiple exchanges between the user and the assistant, while multi-step denotes task completion that requires executing a sequence of dependent actions, often through function calls. Together, they reflect the complexity of real-world agentic scenarios where task states are partially observable. However, constructing such data is inherently challenging: it requires generating complex but solvable tasks, maintaining coherent user-agent exchanges, and accurately simulating tool behaviors. A promising direction lies in multi-agent simulations, where multiple LLMs are assigned roles including user, assistant and tool to collaboratively generate full conversations through autoregressive interactions Liu et al. (2025); Prabhakar et al. (2025). While effective at generating natural conversations, these approaches have several drawbacks: 1) They are computationally costly due to extended back-and-forth interactions, where each new turn must be generated one-by-one based on all previous context; 2) Task complexity and dialogue length are implicitly determined by model interactions and are difficult to constrain explicitly, which poses challenges for fine-grained data design; 3) Most critically, since assistant behaviors are generated autoregressively without access to global context, i.e. the overall task and dependencies between steps, it is difficult for the assistant to optimize the overall output structure and ensure consistency at each step. This lack of holistic awareness hinders factual accuracy, tool-use consistency, and task solvability, especially in scenarios requiring long-term planning. As a result, the quality of generated data largely depends on the capability of the LLMs playing the assistant role, often resembling a form of knowledge distillation from larger (assistant) models.
@@
In this paper, we propose ToolACE-MT, a novel framework for constructing multi-turn dialogues involving agentic tool-use behaviors, inspired by Non-Autoregressive Translation (NAT) and masked diffusion language models Gu et al. (2018); Sahoo et al. (2024), which have been shown to be more efficient in language generation. Unlike traditional autoregressive multi-agent simulations (MAS), ToolACE-MT generates full conversational trajectories through a non-autoregressive pipeline consisting of three stages (see Figure 1): 1) Coarse-Grained Initialization: A coarse but structurally complete dialogue skeleton is generated, specifying overall tasks and action trajectory. 2) Iterative Refinement: Through carefully designed mask-and-fill procedures, the skeleton is progressively enriched with complexity injection and improved with reasonability refinement. 3) Offline Verification: Rule- and model-based checks are applied, filtering out invalid samples.
@@
Exerimental results on several agentic multi-turn benchmarks, including BFCL (Berkeley Function Calling Leaderboard) Yan et al. (2024), -Bench Yao et al. (2025) and ACEBench Chen et al. (2025), show that models trained with ToolACE-MT generated data outperform those with autoregressive multi-agent simulation (MAS). Data analysis and ablation studies confirm the efficiency and effectiveness of our generation pipeline, and more experiments show the generalizability to different backbones.
@@
2 Related Work
Agentic Data Synthesis. LLM agents equipped with external tools have shown realistic problem-solving capabilities Qin et al. (2024); Gou et al. (2024); Lu et al. (2025). As current LLMs still face challenges with complex tasks Mialon et al. (2024); Yao et al. (2025), learning from synthesized agentic data offers a promising direction. Early work focuses on single-turn synthesis, where the agent receives a one-time query and responds accordingly Patil et al. (2023); Zeng et al. (2023); Qin et al. (2024); Liu et al. (2024). However, real-world tasks often involve multi-turn, multi-step interactions, prompting recent studies to explore conversational data synthesis via multi-agent simulations Tang et al. (2023); Liu et al. (2025); Wang et al. (2025). Closest to our approach is Prabhakar et al. (2025), which adopts a two-stage synthesis process. While their first stage resembles ours which generates task configurations and ground-truth answers, they still rely on multi-agent simulations in the second stage to collect full interaction trajectories.
Agentic Model Training. Fine-tuning on synthesized data remains central to agentic model training Qin et al. (2024); Liu et al. (2025); Prabhakar et al. (2025). With reinforcement learning (RL) proving effective in enhancing LLM reasoning Shao et al. (2024); Guo et al. (2025), agentic RL has emerged as a promising alternative for developing agentic capabilities in recent research Feng et al. (2025); Qian et al. (2025); Zhang et al. (2025); Jin et al. (2025). While agentic RL may reduce the complexity of data synthesis by enabling learning from sparse or indirect supervision, high-quality synthesized data remains essential to guide and stabilize training.
Non-Autoregressive Generation. To overcome the inefficiencies and quality limitations in certain scenarios of traditional autoregressive generation, where tokens are produced one-by-one, non-autoregressive approaches have been proposed Gu et al. (2018); Xiao et al. (2023). These methods include CTC-based objectives Libovický and Helcl (2018), iterative refinement with Mask-Predict Ghazvininejad et al. (2019), and insertion and deletion strategies Gu et al. (2019). Inspired by this line of work, we extend the non-autoregressive paradigm to the turn level, enabling more efficient and coherent agentic dialogue data synthesis.
3 Method
3.1 Problem Formulation
@@
This interaction results in a sequence of alternating observations and actions as a trajectory, , where is the user’s initial message and is the tool output or user reply following . constitutes the target for data generation, i.e., multi-turn conversational data.
3.2 Overview
To construct high-quality conversational sequences , prior work applies multi-agent simulations Liu et al. (2025); Prabhakar et al. (2025), where LLMs simulate the user (producing observations after assistant responses), the tools (producing outputs after function calls, can be either actual tools or simulated ones), and the assistant (producing actions). While this approach is shown to work, it is costly, hard to verify and complexity control. We propose a more efficient and controllable method: Non-Autoregressive Iterative Generation (ToolACE-MT), inspired by non-autoregressive translation (NAT) and diffusion models Gu et al. (2018); Sahoo et al. (2024).
Our pipeline consists of three stages: 1) Initialization 2) Iterative Refinement, and 3) Offline Verification. Figure 2 illustrates the overall workflow. We describe each stage in detail below.
3.3 Coarse-Grained Initialization
@@
3.3.1 Task Initialization.
We begin by sampling or specifying a candidate tool list from a predefined tool pool Liu et al. (2025); Wang et al. (2025). The overall task is then generated with the following components: 1) a set of subtasks (with predefined per instance), 2) the required tools for each subtask, 3) and the number of steps required for tool usage for each subtask.
This step serves as high-level planning, outlining the overall trajectory without finalizing all the details. To ensure coverage across domains and promote task diversity, we curate both actual and simulated tools from prior work Qin et al. (2024); Liu et al. (2025) and handcraft initial task examples. The examples will be further enriched during the data generation process.
3.3.2 Trajectory Initialization.
@@
3.4 Iterative Refinement
In this stage, we enhance the initial trajectory through multiple refinement passes, improving both complexity and semantic coherence. Inspired by Masked-Predict Ghazvininejad et al. (2019), we iteratively apply mask-and-extend/fill to progressively improve the trajectory (see Figure 3).
3.4.1 Complexity Injection.
@@ 3.4 Iterative Refinement
Clarification: user gives incomplete information following by assistant’s clarification question.
Tool awareness: assistant recognizes unsupported tasks and user updates the tool list.
Error simulation: tool call fails, resulting assistant reflecting and adjusting actions.
@@
3.5 Offline Verification
Given the extensive use of LLMs in the aforementioned stages, hallucination remains a critical issue, particularly in long multi-turn dialogues involving large tool lists Liu et al. (2023). To address this, we conduct offline verification on the refined trajectories using a hybrid approach that combines rule-based and model-based methods Liu et al. (2024; 2025).
@@ 3.5 Offline Verification
4 Experiments
4.1 Experimental Setup
Dataset Construction. We construct in total 8000 training instances using ToolACE-MT for experiments. For comparison, we also construct 8000 instances with multi-agent simulation (MAS) method introduced in Wang et al. (2025). For fair comparison, we leverage the same candidate tool pool and LLM (GPT-4o-2024-11-20) for generation, the same offline verification is also applied.
Models. We use LLaMA3.1-8B-Instruct AI@Meta (2024) as the base model in our main experiments. Other models, including Qwen2.5-Instruct-series Yang et al. (2024) (0.5B, 1.5B, 3B and 7B) and Qwen3-8B Yang et al. (2025), are also tested to validate the generalizability of our method. To verify the effectiveness of our proposed three stages, we also train models with data without offline verification and iterative refinement for ablation study.
Benchmarks and Evaluation. We conduct experiments on several representative benchmarks targeting on the multi-turn capabilities of tool-augmented LLMs, including the Berkeley Function Call Leaderboard (BFCL-v3) Yan et al. (2024), -Bench Yao et al. (2025), and ACEBench Chen et al. (2025). As we focus on the realistic multi-turn capabilities, we mainly present and analyze results on the categories related to Multi-Turn categories. Results in single turn are also listed (for BFCL, while those for ACEBench listed in Appendix) to show the robustness.
Training Details. Given resource constraints, we employ the parameter-efficient fine-tuning method LoRA Hu et al. (2022) for model training. All model modules are configured for LoRA fine-tuning, with a rank of 16 and an alpha value of 32. Training is performed with a global batch size of 64 and a learning rate of , following a cosine learning rate schedule with a warmup ratio of 0.1.
| Method | Statistics | Quality | |||||||
| Tool # /D | Turn #/D | UT #/D | CT #/D | Call #/UT | SS | EnR | H | Dist-3 | |
| MAS | 8.8 | 28.0 | 5.8 | 8.7 | 2.3 | 65.23 | 43.60 | 7.92 | 0.319 |
| ToolACE-MT | 8.4 | 23.7 | 3.4 | 8.5 | 3.7 | 68.34 | 50.71 | 9.28 | 0.357 |
4.2 Data Quality Assessment
For clarity, Table 1 reports the core data statistics and quality assessment results comparing MAS and ToolACE-MT. In terms of data statistics, ToolACE-MT produces fewer user turns per dialogue but more tool calls per turn. This pattern indicates that we achieve tasks with fewer conversational steps while maintaining a stronger focus on tool invocation, reflecting improved efficiency in multi-step task completion. For quality assessment, following Wang et al. (2025), we evaluate coherence and diversity. Coherence measures the logical consistency and contextual alignment across turns, while diversity captures the variety of linguistic expressions and interaction patterns. The results show that our method yields dialogues with higher coherence and greater diversity.
4.3 Main Results
4.3.1 Results on BFCL
@@
Table 2 shows the results on BFCL-v3. The results demonstrate that ToolACE-MT significantly improves multi-turn function calling accuracy, outperforming strong open-source and even some proprietary models (e.g. Gemini-2.5-Pro). Specifically, ToolACE-MT achieves a multi-turn accuracy, a absolute improvement over the base model Llama3.1-8B-Inst (), and even higher than models with larger sizes like Llama3.1-70B () and DeepSeek-V3 (). Compared to the Multi-Agent Simulation (MAS) (), ToolACE-MT also achieves consistently better results across all multi-turn subcategories. These findings highlight the effectiveness of our ToolACE-MT framework in constructing coherent, contextually grounded dialogues with accurate tool usage.
@@
4.3.2 Results on More Benchmarks
| Models | MT | EA | PA |
@@ 4.3.2 Results on More Benchmarks
| Models | Retail | Airline | Avg. |
@@
The results on -Bench (shown in Table 4) show a consistent trend, with ToolACE-MT outperforming MAS baseline and the ablation models. Interestingly, the base model Llama3.1-8B-Inst obtains a higher score of in the Airline domain, surpassing all trained models. This counterintuitive outcome can be attributed to a known evaluation limitation in -Bench Zhu et al. (2025): several instances define empty actions as the correct responses, assessing the assistant’s ability to recognize unsolvable user requirements. When a model lacks sufficient capability and consistently fails to produce valid function calls, it may coincidentally align with these empty actions and receive positive rewards, despite not demonstrating actual understanding. However, this phenomenon does not persist after training, which ultimately leads to lower evaluation scores.
4.4 Data Efficiency
| Method | Cost | Quality | Performance | |||
| API Call times | Input (#) | Output (#) | Pricing (USD) | |||
| MAS with GPT-4o | 275k | 652,145k | 10,712k | 1,737 | 61.1 | 64.17 |
| ToolACE-MT with GPT-4o | 188k | 478,115k | 18,480k | 1,380 | 72.3 | 65.41 |
| ToolACE-MT with GPT-4o-mini | 394k | 855,391k | 33,062k | 148 | 48.7 | 60.13 |